Circular No. 6880

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ar IN NGC 2916 J.-Y. Wei, Beijing Astronomical Observatory (BAO), write: “A low-

resolution spectrum (range 370–850 nm) of SN 1998ar (IAUC 6877) was obtained on Apr. 18 with the BAO 2.16-m telescope. The spectrum shows a blue continuum and strong, broad Hα and Hβ emissions with P-Cyg profiles, from which SN 1998ar is classified as a type-II supernova around its maximum. The recession velocity of the galactic disk near SN 1998ar is 3750 km/s, as deduced from the narrow Hα emission. The expansion velocity derived from the Hα P-Cyg profile is ∼ 8000 km/s. There are possible Hα and Hβ absorptions from the host galaxy, but a spectrum with higher S/N ratio is needed to confirm these.”

SUPERNOVA 1998aq IN NGC 3982 P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that CCD images obtained by K. Stanek with the Fred L. Whipple Observatory’s 1.2-m telescope shows that SN 1998aq continues to brighten rapidly. Standard V magnitude estimates relative to the star (V = 13) located 16′′ west and 200′′ south of the supernova are: Apr. 17.14 UT, +0.35; 18.14, +0.07; 19.14, −0.14; 20.16, −0.33; this corresponds to a rise of nearly a factor of two in 3 days. Spectra obtained with the 1.5-m Tillinghast telescope by P. Berlind and M. Calkins over the same period shows unusually shallow type-Ia features and a weak Si II absorption with a minimum at 611.7 nm (host galaxy rest frame). The spectrum is very similar to that of SN 1990N, one week before B maximum (Leibundgut et al. 1991, Ap.J. 371, L23).

Visual magnitude estimates: Apr. 15.91 UT, 14.2 (K. Hornoch, Lelekovice, Czech Republic); 15.93, 14.0 (G. Hurst, Basingstoke, England); 15.96, 14.1 (J. Shanklin, Cambridge, England); 17.20, 13.7 (C. E. Spratt, Victoria, BC); 17.83, 13.1 (M. Plsek, Lelekovice, Czech Republic); 17.98, 12.8 (Shanklin); 19.89, 12.7 (P. Schmeer, Bischmisheim, Germany).

AX J0051–73.3 = RX J0050.7–7316 P. C. Schmidtke and A. P. Cowley, Arizona State University, report:

“The optical counterpart of RX J0050.7–7316 (IAUC 6853, 6854, 6860) was identified with a variable peculiar Be star in the Small Magellanic Cloud by Cowley et al. (1997, PASP 109, 21). Its optical position is α = 0^{h}50^{m}44^{s}.8, δ = −73^{o}16′06′′ (equinox J2000.0), with mean values of V = 15.4, B − V = −0.04, and U − B = −0.95 (epoch Nov. 1994).”

1998 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green